OpenStudy (anonymous):

What percent of 12 is 9?

OpenStudy (anonymous):

3/4 *100 = 75 per cent.

OpenStudy (sw050399):

You could think of it as @Anonymous1921 put. Or you could think of it as 9/12 = x/100 since percent is out of 100. Solve for x (cross-multiply).

OpenStudy (anonymous):

\[\frac{ x }{ 100 } \times 12 = 9\]

OpenStudy (anonymous):

find x.

OpenStudy (anonymous):

thnks